# 0000-0002-5763-8174

### v1

@@ 2.4 Dataset
Following TGAS cao2022optimized, we define an $R=1$ acquisition as 50 groups (including the two dummy groups), corresponding to a total scan time of 6 min 38 s. We considered undersampling factors of $R=3,6,$ and $12$, yielding scan times of 2 min 23 s, 1 min 19 s, and 47 s, respectively. For reporting, we use the raw MRF acquisition time (excluding the fixed dummy-group overhead), which we denote as 6 min, 2 min, 1 min, and 30 s for convenience. Practical strategies to bypass the need for these dummy groups are discussed in Section 4.1. All scans were coil-compressed to 10 virtual coils, and coil sensitivity maps were estimated using ESPIRiT uecker2014espirit.
The dataset comprised 44 in vivo subjects scanned on a 3T GE system, with 36 used for training, 4 for validation, and 4 for testing. To assess cross-vendor generalization, we additionally evaluated our method on a subject scanned on a 3T Siemens system. For all subjects, motion was corrected in k-space using navigator data cao2025three. The $R=1$ acquisition was reconstructed using subspace reconstruction with light LLR regularization via FISTA and served as the training target. Quantitative maps were subsequently estimated using a $B_{1}$-corrected dictionary to mitigate systematic $T_{2}$ bias sled2000correction.
To improve numerical stability in SPUR-iG, which performs joint denoising across subspace coefficient maps, we do not directly use the raw SVD basis $\bm{\Phi}$} derived from the MRF signal dictionary, as later components contain substantially less energy. Instead, we apply basis balancing schauman2025deep by performing a discrete Fourier transform along the subspace dimension, $\tilde{{\bm{\Phi}}}={\bf F}_{k}({\bm{\Phi}})$, where ${\bf F}_{k}$ denotes a unitary DFT of length $k$. Since ${\bf F}_{k}$ is orthonormal, this transformation is invertible and leaves the reconstruction solution unchanged, while redistributing energy more evenly across coefficients and thereby stabilizing the training of the learned denoisers. For fair comparison, the balanced basis was also used for all baseline methods.
Finally, we synthesized MPRAGE images from the $R=1$ quantitative maps using Bloch simulations and used them to obtain segmentation masks with FreeSurfer fischl2012freesurfer. These masks were then used to apply region-specific loss weighting during model training.
All human data were acquired with informed consent under protocols approved by the local institutional review board.
@@ 2.5 Evaluation
    LLR: Subspace reconstruction with FISTA and LLR regularization, where the regularization weight was tuned on the validation set separately for each scan duration to minimize $T_{1}$$T_{2}$ error in WM and GM. The number of iterations was set to $40$, ensuring convergence of the loss.
@@
    Hybrid 2D/3D: Inspired by the state-of-the-art SOTIP zou2025improved, this baseline alternates 3D DC updates with axial slice-wise 2D denoising. For a fair comparison, we train this model using the same progressive procedure. Unlike SPUR-iG, the denoiser here is a 2D UNet applied independently to each slice at every unroll iteration, with the 3D volume subsequently formed by stacking the denoised slices. To control for model capacity, we set the initial number of UNet filters such that both this model and SPUR-iG saturated the memory of a single NVIDIA A6000 GPU (48 GB) during fine-tuning.
Separate models were trained for each scan duration. For the unrolled models, we used six unroll iterations. Additional implementation details are provided in Appendix A.
@@ 3 Results
Figure 2:  Example reconstructions of the first balanced subspace coefficient (a) and third unbalanced coefficient (b) for two test subjects. The top row shows the 6 min reference (left) and ${\bf\cal A}^{H}{\bf b}$ for the 30 s and 2 min acquisitions (middle and right) as initialization references. Subsequent rows correspond to reconstructions at 2 min, 1 min, and 30 s, with columns comparing our method, the hybrid 2D/3D variant and LLR. PSNR and SSIM are reported in the bottom right of each image. Error maps of the top-right quadrant (magnified $\times 10$) are overlaid to illustrate error levels.
@@ 3.1 Coefficient Reconstruction
Table. 1 shows the PSNR and SSIM for the five balanced subspace coefficients across scan durations and reconstruction methods. SPUR-iG consistently achieves higher PSNR and SSIM compared to both LLR and the hybrid 2D/3D method, with performance gaps widening as scan duration decreases. Statistical testing confirms that our method outperforms both baselines with $p<0.001$ across all metrics and subspace dimensions.
Qualitative results are shown in Fig. 2 for the first balanced and third unbalanced subspace coefficients. These were chosen to display since balanced coefficients generally display flatter contrast, whereas unbalanced coefficients, particularly the third, exhibit higher tissue contrast and reveal finer anatomical details. As scan time decreases, all methods exhibit increased noise, but our approach shows the smallest degradation. For the 30 s scans, both LLR and the hybrid 2D/3D variant produce noisy and over-smoothed reconstructions, whereas our method better preserves fine structure and anatomical details.
@@ 3.2 Quantitative Reconstruction
Table. 2 shows the average relative quantitative $T_{1}$ and $T_{2}$ errors inside the brain (excluding CSF), defined as $\lvert T^{\text{GT}}-\widehat{T}\rvert/T^{\text{GT}}$ for $T\in\{T_{1},T_{2}\}$, where $\widehat{T}$ denotes the estimated parameter. Across all scan durations, our method achieves lower errors than both baselines. As expected, accuracy decreases with shorter acquisitions, but degradation is less pronounced for our method. Notably, for $T_{1}$, our 30 s scan reconstructions achieved lower average error than LLR reconstructions from a 2 min scan. For $T_{2}$, our approach also provides consistent improvements. Statistical testing confirms that our method outperforms both baselines with $p<0.001$ for $T_{1}$ and $T_{2}$ across all scan durations.
Qualitative results are presented in Fig. 3. For 2 min scans, all methods produce reasonable maps, with ours showing the lowest residual errors. At 30 s, the baseline methods yield overly smoothed maps that lose fine anatomical details, whereas SPUR-iG preserves sharper structures and reduces error.
@@ 3.3 Method Ablation
Fig. 6 shows PSNR and SSIM for the different training variants on 1 min scans (trends are similar for other durations). As can be seen, final fine-tuning provides additional improvements on top of GLEAM, while GLEAM training without pretraining performs poorly. Even when training GLEAM three times longer, performance does not reach that of GLEAM with pretraining.
Iteration conditioning improves performance, and when conditioning is removed, weight sharing decreases performance. Finally, Fig. 7 reports PSNR and SSIM for models trained with a different number of unroll steps, showing that performance improves monotonically and converges around six unroll steps.
@@ 3.4 Out-of-distribution generalization
To assess sensitivity to scanner vendor, we acquired data from a single subject on a Siemens 3T system using the same sequence and reconstructed it using the exact same model and setup (which was trained exclusively on GE data). Figures 4 and 5 show the corresponding subspace reconstructions and quantitative maps, which exhibit trends similar to those observed on the in-distribution test set.
@@
Figure 4:  Reconstructions of the first balanced subspace coefficient (a) and the third unbalanced coefficient (b) for two slices from the same subject, scanned on an out-of-distribution scanner vendor. The layout follows Fig. 2.
@@
Figure 5:  Qualitative examples of reconstructed $T_{1}$ (a) and $T_{2}$ (b) maps from an out-of-distribution vendor scan. The layout follows Fig. 3.
@@ 3.5 Reconstruction Time
Fig. 8 summarizes reconstruction speedups relative to FISTA LLR. Our method achieves up to $\times 111$ speedup for the 2 min acquisition and $\times 51$ for the 30 s acquisition. Compared to iGROG FISTA LLR, the speedups are $\times 15$ and $\times 18$ for the 2 min and 30 s cases, respectively. Notably, reconstruction completes in less than 15 s for all acquisition durations. We assume iGROG implicit kernel training is performed during MRF acquisition after calibration data is obtained and therefore does not add to reconstruction time. Speedup decreases for shorter scans because the DC cost scales with the size of the acquired k-space. The hybrid 2D/3D variant is marginally faster than the 3D model (by less than 1 s), but both provide comparable computational efficiency.
Figure 8:  Relative reconstruction speedup compared to FISTA LLR with 40 iterations. The chart compares iGROG with FISTA LLR (40 iterations), our method with six unroll steps, and the 2D unrolled variant with six unroll steps. For methods using iGROG, reconstruction times include gridding but exclude kernel training (which takes less than 25 s), which can be performed in parallel to acquisition. For reference, the reconstruction times (in seconds) for each method and scan duration, from left to right, are: 1758.6, 241.7, 15.8, 15.0, 0 (2 min); 986, 225.3, 13.1, 12.4 (1 min); and 601.7, 217.0, 11.8, 11.0 (30 s). All experiments are performed on a single A6000 GPU.
@@ 4.1 Quantitative Estimation
With aggressive undersampling, LLR requires stronger regularization to suppress noise, which introduces systematic bias and produces overly smooth maps (Fig. 3, rightmost column). This is most evident at tissue boundaries, where the low-rank assumption does not hold. Although we shift patches during LLR reconstruction to reduce blocking artifacts, the patch-based constraint still couples signals from different tissue types, leading to biased values at boundaries (e.g., elevated tissue values for WM and reduced values for GM or CSF).
Our learned prior mitigates this effect by replacing patch-based low-rank constraints with a UNet denoiser. Unlike LLR, the UNet produces voxel-wise predictions informed by convolutional receptive fields, allowing it to incorporate spatial context while preserving local tissue differences. In practice, this enables the network to recognize and respect boundaries rather than enforcing a shared representation across them. Residual averaging effects remain, especially at shorter scan durations, but they are substantially smaller. The 3D model further reduces this bias by leveraging volumetric context. Most of the remaining error appears in CSF, which is expected since this particular MRF sequence is primarily designed to optimize sensitivity for WM and GM rather than CSF cao2022optimized.
Although our training objective is not explicitly designed to optimize tissue quantification, but rather to improve subspace coefficient reconstruction, Table. 2 shows that our method improves quantitative performance, highlighting the benefit of the learned prior. Future work could integrate quantitative parameter estimation directly into the pipeline and jointly optimize for estimation accuracy zou2025improved, or to optimize the learned prior for other downstream applications, such as multi-compartment modeling.
@@ 4.2 Coefficient Maps
Coefficient maps preserve richer temporal information than $T_{1}/T_{2}$ maps and are important for emerging downstream applications such as multi-compartment modeling and clinical contrast synthesis. As shown in Table. 1, our method consistently outperforms both baselines in PSNR and SSIM, with performance gaps widening as scan duration decreases. The improvements come from two factors: the learned prior, which captures more accurate image statistics than hand-crafted regularization, and the 3D UNet architecture, which leverages volumetric correlations across slices for stronger regularization. The proposed training strategy enables training and deployment of such a model, supporting a more complex and accurate representation of image statistics and, consequently, improved reconstructions.
Future work could further improve performance by using optimized spatio-temporal k-space sampling trajectories zou2025improved, which may be especially valuable for very short acquisitions such as 30 s.
@@ 4.3 Method Design
Fig. 6 evaluates the impact of different components of our framework.
Pretraining. Pretraining is a key element of our approach. By avoiding unrolling and DC, it enables efficient training with extensive data augmentation, which improves generalization and robustness. Without pretraining, GLEAM performs poorly, and even tripling the training time is insufficient to reach comparable performance.
Final fine-tuning. Fine-tuning with full unrolling provides a modest but consistent improvement by directly optimizing the final reconstruction loss.
Iteration conditioning. Conditioning on the unroll iteration index improves performance over unconditional variants. It simplifies the denoiser’s task by explicitly providing the expected artifact and noise-level distribution, rather than requiring the model to infer these solely from the input. In unconditional training, not sharing weights performs better than sharing, since the model no longer has to infer denoising strength from the input. However, both remain inferior to the conditioned version, which amortizes learning across noise regimes, effectively enlarging the dataset. The cost of conditioning is negligible, adding less than 1% to the model parameters.
Efficiency. These design choices enhance performance while achieving low runtime. The learned denoiser is effective at modeling complex distributions, and its inference is efficient, especially on modern GPUs, leading to reduced reconstruction times.
Although the training set consists exclusively of data acquired on a 3T GE system, Figs. 4 and 5 demonstrate that the method generalizes to an out-of-distribution vendor without retraining or fine-tuning, indicating strong cross-vendor robustness. While a modest performance degradation is observed, we expect this gap could be mitigated by incorporating a small amount of out-of-distribution data during training or fine-tuning.
@@ 5 Conclusions
We propose a fully 3D unrolled MRF reconstruction framework with a staged training strategy that enables large-scale training within practical compute limits. By combining efficient iGROG-based DC with fast 3D UNet-based inference, the method achieves rapid reconstruction while consistently improving quality, supporting acquisitions as short as 30 s and preserving fidelity in both subspace coefficients and quantitative maps

### v2

@@ 2.4 Dataset
Following the convention from TGAS cao2022optimized, we define an $R=1$ acquisition as 50 groups (including the two dummy groups), corresponding to a total scan time of 6 min 38 s. We considered undersampling factors of $R=3,6,$ and $12$, yielding scan times of 2 min 23 s, 1 min 19 s, and 47 s, respectively. For reporting, we use the raw MRF acquisition time (excluding the fixed dummy-group overhead), which we round and denote as 6 min, 2 min, 1 min, and 30 s for convenience. Practical strategies to bypass the need for these dummy groups are discussed in Section 4.1. All scans were coil-compressed to 10 virtual coils, and coil sensitivity maps were estimated using ESPIRiT uecker2014espirit.
The dataset comprised 44 in vivo subjects scanned on a 3T GE SIGNA Premier system, with 36 used for training, 4 for validation, and 4 for testing. To assess cross-vendor generalization, we additionally evaluated our method on a subject scanned on a 3T Siemens MAGNETOM Vida system. For all subjects, motion was corrected in k-space using navigator data cao2025three. The $R=1$ acquisition was reconstructed using subspace reconstruction with light LLR regularization via FISTA and served as the training target. Quantitative maps were subsequently estimated using a $B_{1}$-corrected dictionary to mitigate systematic $T_{2}$ bias sled2000correction.
To improve numerical stability in SPUR-iG, which performs joint denoising across subspace coefficient maps, we do not directly use the raw SVD basis $\bm{\Phi}$} derived from the MRF signal dictionary, as later components contain substantially less energy. Instead, we apply basis balancing schauman2025deep by performing a discrete Fourier transform along the subspace dimension, $\tilde{{\bm{\Phi}}}={\bf F}_{k}({\bm{\Phi}})$, where ${\bf F}_{k}$ denotes a unitary DFT of length $k$. Since ${\bf F}_{k}$ is orthonormal, this transformation is invertible and leaves the reconstruction solution unchanged, while redistributing energy more evenly across coefficients and thereby stabilizing the training of the learned denoisers. For a fair comparison, the balanced basis was used for all baseline methods.
Finally, we synthesized MPRAGE images from the $R=1$ quantitative maps using Bloch simulation and used them to obtain segmentation masks with FreeSurfer fischl2012freesurfer. These masks were then used to apply region-specific loss weighting during model training (additional details in Appendix A.2).
@@ 2.5 Evaluation
    LLR: Subspace reconstruction with FISTA and LLR regularization. The regularization weight was tuned on the validation set for each scan duration to minimize the sum of the mean relative $T_{1}$ and $T_{2}$ errors within WM and GM. The number of iterations was fixed at $40$, ensuring convergence.
@@
    Hybrid 2D/3D: Inspired by the state-of-the-art SOTIP zou2025improved, this baseline alternates 3D DC updates with axial slice-wise 2D denoising. For a fair comparison, we train this model using the same progressive procedure. Unlike SPUR-iG, the denoiser here is a 2D UNet applied independently to each slice at every unroll iteration, with the 3D volume subsequently formed by stacking the denoised slices. To control for model capacity, we increased the number of parameters in the 2D UNet such that both this model and SPUR-iG saturated the memory of a single NVIDIA A6000 GPU (48 GB) during the final fine-tuning stage
Separate models were trained for each scan duration. For the unrolled models, we used six unroll iterations, selected based on ablation studies (Section 3.3). Additional implementation details are provided in Appendix A.
@@ 3 Results
Figure 2:  Example reconstructions of the first balanced subspace coefficient (a) and third unbalanced coefficient (b) for two test subjects. The top row shows the 6 min reference (left) and ${\bf\cal A}^{H}{\bf b}$ for the 30 s and 2 min acquisitions (middle and right) as initialization references. Subsequent rows correspond to reconstructions at 2 min, 1 min, and 30 s, with columns comparing our method, the hybrid 2D/3D variant and LLR. PSNR and SSIM are reported in the bottom right of each image. Error maps of the top-right quadrant (magnified $\times 10$ or $\times 5$) are overlaid to illustrate error levels.
@@ 3.1 Coefficient Reconstruction
Table 1 shows the PSNR and SSIM for the five balanced subspace coefficients across scan durations and reconstruction methods. SPUR-iG consistently achieves higher PSNR and SSIM compared to both LLR and the hybrid 2D/3D method, with performance gaps widening as scan duration decreases. Statistical testing confirms that our method outperforms both baselines with $p<0.001$ across all metrics and subspace dimensions.
Representative qualitative results are shown in Figure 2 for the first balanced and third unbalanced subspace coefficients. These were chosen to display since balanced coefficients generally display flatter contrast, whereas unbalanced coefficients, particularly the third, exhibit higher tissue contrast and reveal finer anatomical details. As scan time decreases, all methods exhibit increased noise, but our approach shows the smallest degradation. For the 30 s scans, both LLR and the hybrid 2D/3D variant produce noisy and over-smoothed reconstructions, whereas our method better preserves fine structure and anatomical details.
@@ 3.2 Quantitative Reconstruction
Table 2 shows the average relative quantitative $T_{1}$ and $T_{2}$ errors inside the brain (excluding CSF), defined as $\lvert T^{\text{GT}}-\widehat{T}\rvert/T^{\text{GT}}$ for $T\in\{T_{1},T_{2}\}$, where $\widehat{T}$ denotes the estimated parameter. Across all scan durations, our method achieves lower errors than both baselines. As expected, accuracy decreases with shorter acquisitions, but degradation is less pronounced for our method. Notably, for $T_{1}$, our reconstructions of the 30 s scans achieved lower average error than LLR reconstructions of 2 min scans. For $T_{2}$, our approach also provides consistent improvements. Statistical testing confirms that our method outperforms both baselines with $p<0.001$ for $T_{1}$ and $T_{2}$ across all scan durations.
Representative qualitative results are presented in Figure 3. For 2 min scans, all methods produce reasonable maps, with ours showing the lowest residual errors. At 30 s, the baseline methods yield overly smoothed maps that lose fine anatomical details, whereas SPUR-iG preserves sharper structures and reduces error.
@@ 3.3 Method Ablation
Figure 4 shows PSNR and SSIM for the different training variants on 1 min scans (trends are similar for other durations). As can be seen, final fine-tuning provides additional improvements on top of GLEAM, while GLEAM training without pretraining performs poorly. Even when training GLEAM three times longer, performance does not reach that of GLEAM with pretraining.
Iteration conditioning improves performance, and when conditioning is removed, weight sharing decreases performance. Finally, Figure 5 reports PSNR and SSIM for models trained with a different number of unroll steps, showing that performance improves monotonically and converges around six unroll steps.
@@ 3.4 Out-of-distribution generalization
To assess sensitivity to scanner vendor, we acquired data from a single subject on a Siemens 3T system using the same sequence and reconstructed it using the exact same model and setup (which was trained exclusively on GE data). Figures 6 and 7 show the corresponding subspace reconstructions and quantitative maps, which exhibit trends similar to those observed on the in-distribution test set.
@@
Figure 6:  Reconstructions of the first balanced subspace coefficient (a) and the third unbalanced coefficient (b) for two slices from the same subject, scanned on an out-of-distribution scanner vendor. The layout follows Figure 2.
@@
Figure 7:  Qualitative examples of reconstructed $T_{1}$ (a) and $T_{2}$ (b) maps from an out-of-distribution vendor scan. The layout follows Figure 3.
@@ 3.5 Reconstruction Time
Figure 8 summarizes reconstruction speedups relative to FISTA LLR. All experiments were performed on a Linux workstation using a single NVIDIA A6000 GPU (48GB memory) with a PyTorch implementation. Our method achieves up to $\times 111$ speedup for the 2 min acquisition and $\times 51$ for the 30 s acquisition. Compared to iGROG FISTA LLR, the speedups are $\times 15$ and $\times 18$ for the 2 min and 30 s cases, respectively. Notably, reconstruction completes in less than 15 s for all acquisition durations. We assume iGROG INR training is performed during MRF acquisition after calibration data are obtained and therefore does not add to reconstruction time, while gridding is included in the reported reconstruction times (kernel training completes in under 25 s and gridding in under 3 s across all acquisitions considered here). Speedup decreases for shorter scans as the DC cost scales with the size of the acquired k-space. The hybrid 2D/3D variant is marginally faster than the 3D model (by less than 1 s), but both provide comparable computational efficiency.
Figure 8:  Relative reconstruction speedup compared to FISTA LLR with 40 iterations. The chart compares iGROG with FISTA LLR (40 iterations), our method with six unroll steps, and the hybrid 2D/3D unrolled variant with six unroll steps. For methods using iGROG, reconstruction times include gridding but exclude kernel training (which takes less than 25 s, and can be performed in parallel to acquisition). For reference, the reconstruction times (in seconds) for each method and scan duration, from left to right, are: 1758.6, 241.7, 15.8, 15.0 (2 min); 986, 225.3, 13.1, 12.4 (1 min); and 601.7, 217.0, 11.8, 11.0 (30 s). All experiments are performed on a single A6000 GPU.
@@ 4.1 Quantitative Estimation
With aggressive undersampling, LLR requires stronger regularization to suppress noise, which introduces systematic bias and produces overly smooth maps (Figure 3, rightmost column). This is most evident at tissue boundaries, where the low-rank assumption does not hold. Although we jitter patch locations across iterations during LLR reconstruction to reduce blocking artifacts, the patch-based constraint still couples signals from different tissue types, leading to biased values at boundaries (e.g., elevated tissue values for WM and reduced values for GM or CSF).
Our learned prior mitigates this effect by replacing patch-based low-rank constraints with a UNet denoiser. Unlike LLR, the UNet produces voxel-wise predictions informed by convolutional receptive fields, allowing it to incorporate spatial context while preserving local tissue differences. In practice, this enables the network to recognize and respect boundaries rather than enforcing a shared representation across them. Residual averaging effects remain, especially at shorter scan durations, but they are substantially smaller. The 3D model further reduces this bias by leveraging volumetric context. Most of the remaining error appears in CSF, which is expected since this particular MRF sequence is primarily designed to optimize sensitivity for WM/GM cao2022optimized.
Although our training objective is not explicitly designed to optimize tissue quantification, but rather to improve subspace coefficient reconstruction, Table 2 shows that our method improves quantitative performance, highlighting the benefit of the learned prior. Future work could integrate quantitative parameter estimation directly into the pipeline and jointly optimize for estimation accuracy, or to optimize the learned prior for other downstream applications, such as multi-compartment modeling.
@@ 4.2 Coefficient Maps
Coefficient maps preserve richer temporal information than $T_{1}/T_{2}$ maps and are important for emerging downstream applications such as multi-compartment modeling and clinical contrast synthesis. As shown in Table 1, our method consistently outperforms the baselines in PSNR and SSIM, with performance gaps widening as scan duration decreases. The improvements come from two factors: the learned prior, which captures more accurate image statistics than hand-crafted regularization, and the 3D UNet architecture, which leverages volumetric correlations across slices for stronger regularization. The proposed training strategy enables training and deployment of such a model, supporting a more complex and accurate representation of image statistics and, consequently, improved reconstructions.
The MRF scans used in this work utilize an optimized spatio-temporal k-space sampling strategy based on the TGAS approach cao2022optimized. Future work could further improve performance by refining sampling trajectories within the reconstruction training process zou2025improved, which may be particularly valuable for very short acquisitions such as 30 s.
@@ 4.3 Method Design
Figure 4 evaluates the impact of different components of our framework:
  1. 1.
    Pretraining is a key element of our approach. By avoiding unrolling and DC, it enables efficient training with extensive data augmentation, which improves generalization and robustness. Without pretraining, GLEAM performs poorly, and even tripling the training time is insufficient to reach comparable performance.
  2. 2.
    Fine-tuning with full unrolling provides a modest but consistent improvement by directly optimizing the final reconstruction loss.
  3. 3.
    Conditioning on the unroll iteration index improves performance over unconditional variants. It simplifies the denoiser’s task by explicitly providing the expected artifact and noise-level distribution, rather than requiring the model to infer these solely from the input. In unconditional training, not sharing weights performs better than sharing, since the model no longer has to infer denoising strength from the input. However, both remain inferior to the conditioned version, which amortizes learning across noise regimes, effectively enlarging the dataset. The cost of conditioning is negligible, adding less than 1% to the model parameters.
These design choices enhance performance while achieving low runtime. The learned denoiser is effective at modeling complex distributions, and its inference is efficient, especially on modern GPUs, leading to reduced reconstruction times.
Although the training set consists exclusively of data acquired on a 3T GE system, Figs. 6 and 7 demonstrate that the method generalizes to an out-of-distribution vendor without retraining or fine-tuning, indicating cross-vendor robustness. While some performance degradation is observed, we expect this gap could be mitigated by incorporating a small amount of out-of-distribution data during training or fine-tuning.
@@ 5 Conclusions
We propose a fully 3D unrolled MRF reconstruction framework with a staged training strategy that enables large-scale training within practical compute limits. By combining efficient iGROG-based DC with fast 3D UNet-based inference, the method achieves rapid reconstruction while consistently improving quality, supporting acquisitions as short as 30 s and preserving fidelity in both subspace coefficients and quantitative maps.